14) Express the integrand as a sum of partial fractions and evaluate the integral

\int \frac{x^{2}+4x+5}{(x-1)(x^{2}+4)}dx
Answer #1

n² +4x+5 (n-1) (2+4) using Partial fraction decomposition +4215 Bxt 6 (2-1) (2244) - A & But 6 x²+4) N-1 = x²+42 +5 (0-1) (A=2 in eg. (11 put 442-(=5 (8-5 (C=3] Hence +4215 (2-)) (x²+4) x+3 2244 now a dat & go to an dla pt dn t f 7 - 12 de 2 dn 2-1